# Generative Diffusion Models for High Dimensional Channel Estimation

### v2

@@ I Introduction
Notations: For any matrix $\mathbf{A}$, $\mathbf{A}^{T}$, $\mathbf{A}^{\rm H}$, and $\mathbf{A}^{-1}$ represent the transpose, conjugate transpose, and inverse of $\mathbf{A}$, respectively. Also, $\mathbf{0}$ is a zero vector, $\mathbf{I}$ is an identity matrix, $\|\cdot\|_{2}$ denotes the $l_{2}$-norm, $\mathrm{vec}(\cdot)$ denotes the vectorization operation, $\otimes$ denotes the Kronecker product, $\mathbb{E}[\cdot]$ denotes the expectation operator, and $\mathbb{I}(\cdot)$ represents an indicator function that returns value one only when the event as the argument is true and zero otherwise. Moreover, $\mathrm{Re}(\cdot)$ and $\mathrm{Im}(\cdot)$ extract the real and imaginary parts of a complex value, respectively. $\mathcal{N}(z;\mu,\sigma^{2})$ represents a Gaussian random variable $z$ with mean $\mu$ and variance $\sigma^{2}$. Finally, $\mathbb{C}$ and $\mathbb{R}$ denote the sets of complex and real numbers, respectively, $[N]$ is the set of nonnegative integers up to $N$, and $\mathcal{U}([N])$ denotes a discrete uniform distribution within $[N]$.
@@ II-A MIMO Channel Estimation
where $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{N_{\rm p}}]$ and $\mathbf{N}=[\mathbf{n}_{1},\ldots,\mathbf{n}_{N_{\rm p}}]$ contain the stacked pilot symbols and noise vectors, respectively. For simplicity, the pilot symbols are supposed to be equiprobably drawn from the quadrature phase shift keying (QPSK) lattice, i.e., {±$\{\pm 1/\sqrt{2}\pm{\rm j}/\sqrt{2}\}$, fulfilling the unit power constraint.
@@ II-B Denoising Diffusion Generative Models
where $\mathbf{h}_{t}$ is the latent variable at time step $t$, and $10<\beta_{1}<\beta_{2}<\ldots<\beta_{T}<$ 1 denote the prescribed noise schedule. Defining $\alpha_{t}\triangleq 1-\beta_{t}$ and $\bar{\alpha}_{t}\triangleq\prod_{i=1}^{t}\alpha_{i}$ and using the reparameterization trick, it is easy to derive that
@@ III-A2 Derivation of the Posterior Inference Algorithm
0:  ${\bf{A}}$, ${\bf{y}}$, $\sigma_{n}^{2}$, pre-trained denoising network $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$, noise schedule $\{\beta_{t}\}_{t=1}^{T}$, gradient scale $s$.
@@ III-B DM-Based Quantized Channel Estimation
For ADCs with $b$ quantization bits, the quantizer maps the input into a countable set of codewords, =$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{2^{b}-1}\}$. The codewords correspond to intervals split by the quantization thresholds $\mathcal{T}=\{(r_{k}^{\rm low},r_{k}^{\rm up})\mid k\in[2^{b}]\}$. For example, if $Q(z)=\bar{z}\in\mathcal{R}$, the input $z$ would satisfy $\bar{z}^{\rm low}\leq z<\bar{z}^{\rm up}$, where $\bar{z}^{\rm low}$ and $\bar{z}^{\rm up}$ represent the lower and upper thresholds with respect to the codeword $\bar{z}$, respectively. Considering a uniform mid-rise quantizer [^5] The proposed channel estimation method is not confined to this choice and can be readily extended to more complex non-uniform quantizers., we have $r_{k}=(2k-2^{b}-1)\frac{\Delta}{2},\;k\in[2^{b}]$,
@@
$\bar{\mathbf{y}}=Q(\frac{1}{\sqrt{\bar{\alpha}_{t}}}\mathbf{Ah}_{t}+\tilde{\mathbf{n}}),$ (26)
@@ III-C Learning from Noisy Channel Realizations
Inspired by [43], we develop a training strategy for the DM when only a noisy channel dataset ¯=$\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}$ is available by coupling the SURE denoising into the training of DM. To begin, we provide a brief introduction to the SURE denoising.
@@
$\mathcal{L}$ (${}_{\mathrm{SURE},\sigma_{w}}(\boldsymbol{\theta})$
@@
To avoid the high-complexity $N$-times partial derivates involved in this definition, we follow the well-established practice [46] to simplify computation by employing a Monte Carlo approximation as follows
@@ IV-A Experimental Details
We adopt the QuaDRiGa toolbox [49] as the channel simulator to generate training, testing, and validation data. We consider an urban macro-cell line-of-sight (LOS) scenario according to the 3rd generation partnership project (3GPP) 38.901 technical report [52]. The carrier frequency is set to 40 GHz. ULAs with “3GPP-3D” antennas featuring half-wavelength spacing are utilized at both the base station (BS) and the multi-antenna user terminal (UT). Downlink transmission is assumed, where the BS’s antenna array is placed at a height of 25 m. The UT randomly resides in a $120^{\circ}$ cell sector with a radius of 500 m and a guard distance of 35 m to the BS. The number of antennas at the BS and UT is set as $N_{\rm t}=64$ and $N_{\rm r}=16$, respectively. A total of $D=100,000$ channel samples are generated as the training dataset, each with different UT locations to enhance diversity. Two distinct setups are considered, i.e., noiseless and noisy datasets. For the noiseless case (Sections IV-B and IV-C), a simulated dataset =$\mathcal{H}=\{\mathbf{h}^{(i)}\}_{i=1}^{D}$ containing ground truth samples generated via QuaDRiGa is employed; for the noisy case (Section IV-E), the simulated dataset $\mathcal{H}$ serves as a benchmark for deriving normalized LMMSE estimates as given in (32), constituting a noisy dataset ¯=$\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$. Furthermore, a validation dataset containing 100 channel realizations is utilized for hyperparameter tuning.
@@ IV-E Effectiveness of SURE-DM
Previous subsections assume the availability of ground truth channel samples. Next, we target the case where only noisy channel data is accessible to evaluate the effectiveness of the proposed SURE-DM. We first train the MMSE denoiser for 100 epochs using the SURE loss given the noisy dataset ¯=$\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$, and then train the DM for 500 epochs using the MMSE-denoised samples corresponding to $\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$.
@@ Appendix B Proof of Proposition 1
=$\int$= ∫ $p_{t-1}(\mathbf{h}_{t-1}|\mathbf{y})p_{t}(\mathbf{h}_{t}|\mathbf{h}_{t-1},\mathbf{y})$
@@ Appendix C Proof of Proposition 2
where $p_{\tilde{n}}(\cdot)$ is the probability density function of $\tilde{n}$, with the index $m$ omitted for brevity. Based on the definition of the quantizer $Q$, we have

### v3

@@ I Introduction
Notations: For any matrix $\mathbf{A}$, $\mathbf{A}^{T}$, $\mathbf{A}^{\rm H}$, and $\mathbf{A}^{-1}$ represent the transpose, conjugate transpose, and inverse of $\mathbf{A}$, respectively. Also, $\mathbf{0}$ is a zero vector, $\mathbf{I}$ is an identity matrix, $\|\cdot\|_{2}$ denotes the $l_{2}$-norm, $\mathrm{vec}(\cdot)$ denotes the vectorization operation, $\otimes$ denotes the Kronecker product, $\mathbb{E}[\cdot]$ denotes the expectation operator, and $\mathbb{I}(\cdot)$ represents an indicator function that returns value one only when the event as the argument is true and zero otherwise. Moreover, $\mathrm{Re}(\cdot)$ and $\mathrm{Im}(\cdot)$ extract the real and imaginary parts of a complex value, respectively. $\mathcal{N}(z;\mu,\sigma^{2})$ represents a Gaussian random variable $z$ with mean $\mu$ and variance $\sigma^{2}$. Finally, $\mathbb{C}$ and $\mathbb{R}$ denote the sets of complex and real numbers, respectively, $[N]$ is the set of nonnegative integers up to NN, and $\mathcal{U}([N])$ denotes a discrete uniform distribution within $[N]$.
@@ II-A MIMO Channel Estimation
where $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{N_{\rm p}}]$ and $\mathbf{N}=[\mathbf{n}_{1},\ldots,\mathbf{n}_{N_{\rm p}}]$ contain the stacked pilot symbols and noise vectors, respectively. For simplicity, the pilot symbols are supposed to be equiprobably drawn from the quadrature phase shift keying (QPSK) lattice, i.e., $\{\pm 1/\sqrt{2}\pm{\rm j}/\sqrt{2}\}$, fulfilling the unit power constraint.
@@ II-B Denoising Diffusion Generative Models
where $\mathbf{h}_{t}$ is the latent variable at time step $t$, and $0<\beta_{1}<\beta_{2}<\ldots<\beta_{T}<1$ denote the prescribed noise schedule. Defining $\alpha_{t}\triangleq 1-\beta_{t}$ and $\bar{\alpha}_{t}\triangleq\prod_{i=1}^{t}\alpha_{i}$ and using the reparameterization trick, it is easy to derive that
@@ III-A2 Derivation of the Posterior Inference Algorithm
[…]
@@ III-B DM-Based Quantized Channel Estimation
For ADCs with $b$ quantization bits, the quantizer maps the input into a countable set of codewords, $\mathcal{R}=\{r_{1},r_{2},\ldots,r_{2^{b}-1}\}$. The codewords correspond to intervals split by the quantization thresholds $\mathcal{T}=\{(r_{k}^{\rm low},r_{k}^{\rm up})\mid k\in[2^{b}]\}$. For example, if $Q(z)=\bar{z}\in\mathcal{R}$, the input $z$ would satisfy $\bar{z}^{\rm low}\leq z<\bar{z}^{\rm up}$, where $\bar{z}^{\rm low}$ and $\bar{z}^{\rm up}$ represent the lower and upper thresholds with respect to the codeword $\bar{z}$, respectively. Considering a uniform mid-rise quantizer [^5] The proposed channel estimation method is not confined to this choice and can be readily extended to more complex non-uniform quantizers., we have $r_{k}=(2k-2^{b}-1)\frac{\Delta}{2},\;k\in[2^{b}]$,
@@
[…]
@@ III-C Learning from Noisy Channel Realizations
Inspired by [43], we develop a training strategy for the DM when only a noisy channel dataset $\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}$ is available by coupling the SURE denoising into the training of DM. To begin, we provide a brief introduction to the SURE denoising.
@@
[…]
@@
To avoid the high-complexity NN-times partial derivates involved in this definition, we follow the well-established practice [46] to simplify computation by employing a Monte Carlo approximation as follows
@@ IV-A Experimental Details
We adopt the QuaDRiGa toolbox [49] as the channel simulator to generate training, testing, and validation data. We consider an urban macro-cell line-of-sight (LOS) scenario according to the 3rd generation partnership project (3GPP) 38.901 technical report [52]. The carrier frequency is set to 40 GHz. ULAs with “3GPP-3D” antennas featuring half-wavelength spacing are utilized at both the base station (BS) and the multi-antenna user terminal (UT). Downlink transmission is assumed, where the BS’s antenna array is placed at a height of 25 m. The UT randomly resides in a $120^{\circ}$ cell sector with a radius of 500 m and a guard distance of 35 m to the BS. The number of antennas at the BS and UT is set as $N_{\rm t}=64$ and $N_{\rm r}=16$, respectively. A total of $D=100,000$ channel samples are generated as the training dataset, each with different UT locations to enhance diversity. Two distinct setups are considered, i.e., noiseless and noisy datasets. For the noiseless case (Sections IV-B and IV-C), a simulated dataset $\mathcal{H}=\{\mathbf{h}^{(i)}\}_{i=1}^{D}$ containing ground truth samples generated via QuaDRiGa is employed; for the noisy case (Section IV-E), the simulated dataset $\mathcal{H}$ serves as a benchmark for deriving normalized LMMSE estimates as given in (32), constituting a noisy dataset $\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$. Furthermore, a validation dataset containing 100 channel realizations is utilized for hyperparameter tuning.
@@ IV-E Effectiveness of SURE-DM
Previous subsections assume the availability of ground truth channel samples. Next, we target the case where only noisy channel data is accessible to evaluate the effectiveness of the proposed SURE-DM. We first train the MMSE denoiser for 100 epochs using the SURE loss given the noisy dataset $\bar{\mathcal{H}}=\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$, and then train the DM for 500 epochs using the MMSE-denoised samples corresponding to $\{\bar{\mathbf{h}}^{(i)}\}_{i=1}^{D}$.
@@ Appendix B Proof of Proposition 1
$=\int$ $p_{t-1}(\mathbf{h}_{t-1}|\mathbf{y})p_{t}(\mathbf{h}_{t}|\mathbf{h}_{t-1},\mathbf{y})$
@@ Appendix C Proof of Proposition 2
[…]